OpenStudy (anonymous):

3 = log8 + 3logx. Solve for x.

OpenStudy (anonymous):

I have to solve for x.

OpenStudy (anonymous):

First get the 3 in 3logx inside the log as a power so it will be logx^3 Now use the law : Logx+logy=logxy . And according : 3=log8x^3 Now change it to exponential 10^3=8x^3 1000=8x^3 ---> divide by 8 125=x^3 --->take cubic root X=5 _____________________________________- Any questions ?

OpenStudy (anonymous):

Oh okay, Thank you! :)

OpenStudy (anonymous):

Yw.